the equation of a curve is given as f(x)= sqrt 2x-1, find the equation of the tangent line to the graph of f(x) at x=5


1
Expert's answer
2021-11-16T11:24:10-0500

"y=\sqrt{2x-1}"

"y(5)=\sqrt{10-1}=3"

"y'=\frac{1}{\sqrt{2x-1}}"

"y'(5)=\frac{1}{\sqrt{10-1}}=\frac{1}{3}"

Tangent line: "y-3=\frac{1}{3}(x-5)" or "y=\frac{1}{3}x+\frac{4}{3}"